OpenStudy (anonymous):

A 9.0 kg box is sliding down a smooth flat ramp which makes an angle of 24° with the horizontal. If the coefficient of kinetic friction between the box and the ramp is 0.25, how long will it take the box to slide 2.0 m down the ramp from a standing start?

OpenStudy (anonymous):

First determine the acceleration by summing up the forces in the x-direction, along the ramp, and setting it equal to ma. Using down the ramp as the positive direction: Fx = mgsin(theta) - uk mgcos(theta) = ma uk is the coefficient of kinetic friction With the acceleration, you can use the equations of motion with constant acceleration to determine the time: x2 = x1 + v1t + 1/2 a t^2 where x1 = 0 and v1 = 0

OpenStudy (anonymous):

Excellent! Thank you! :D